Reserve Problems Chapter 11 Section 2 Problem 1 The department of health studied the number of patients who need liver transpReserve Problems Chapter 11 Section 2 Problem 2 The peanut crop was harvested from five fields of various area. The followingReserve Problems Chapter 12 Section 1 Problem 11 An article in Technometrics (1974, Vol. 16, pp. 523-531) considered the foll15 70 20 91 (a) Fit a linear regression model relating the results of the stack loss to the three regressor variables. RoundReserve Problems Chapter 12 Section 1 Problem 1 During a research, the amount of Internet users was measured. Each time three
Answer #1

Solution:

The basic concept of regression models and its parameters and estimates calculation is needed as a prerequisite to under the solution to the given problem. Since a lot of data are there, I have used R program to calculate the necessary requirements, parameter estimation, calculation of model variance & necessary other calculations required. The 4 datasets are named d1, d2, d3 & d4 and all the calculations for each dataset has been done seperately with the models named as 'model1', 'm2', 'm3' & 'm4' respectively. The outputs for each of the dataset are mentioned in the comments under each of the codes done for the seperate datasets. The solution is pasted below:

######Reading 4 seperate datasets######

x1=c(2003:2015)
y1=c(1278,1761,2917,3923,4983,6914,7655,7596,7737,8524,8376,8639,8946)
d1=data.frame(x1,y1)

x2=c(1.91,4.29,3.43,4.01,3.05)
y2=c(7300,15510,13650,19940,13100)
d2=data.frame(x2,y2)

x1_=c(80,80,75,rep(62,5),rep(58,6),rep(50,5),56,70)
x2_=c(rep(27,2),25,24,22,23,24,24,23,18,18,17,18,19,18,18,19,19,rep(20,3))
x3_=c(89,88,90,rep(87,3),rep(93,2),87,80,89,88,82,93,89,86,72,79,80,82,91)
y_=c(42,37,37,28,18,18,19,20,15,14,14,13,11,12,8,7,8,8,9,15,15)
d3=data.frame(x1_,x2_,x3_,y_)

x11=c(1,1,1,3,3,3,5,5,5,7,7,7)
x21=c(20,40,60,20,40,60,20,40,60,20,40,60)
y11=c(1500,1250,180,3100,2600,530,4550,4210,910,6290,5730,1320)
d4=data.frame(x11,x21,y11)

#------------------------------------------------------------------------------------------------------------------------#

#######Basic Regression and estimation of parameters#######


######Start of Data1#######

#Data 1 = d1

#Fit a model

model1=lm(y1~x1,data=d1)
model1
summary(model1)
sigma_sq1=(summary(model1)$sigma)**2
b1=summary(model1)$coef[,1][2]
b1
Y_2018 = -1343102+671.6*2018
Y_2018
Y_2012=-1343102+671.6*2012
res1=8524-Y_2012


#The model is
# Y = -1343102+671.6*X
# The estimate of model variance is 882970.4
# The regression coefficient is 671.6
# The value of Y in 2018 is 12140.23
# The value of Y in 2012 is 8157.2
# The residual in Y in 2012 is 366.8

########End of Data 1######

#----------------------------------------------------------------------------------------------------------------#

######Start of Data2######

#Data 2 = d2

# Fit a model

m2=lm(y2~x2,data=d2)
sigma_sq2=(summary(m2)$sigma)**2
b2=summary(m2)$coef[,1][2]
y_4th=-667.1+4364*4.01
res2_4th_val_x=19940-y_4th
y_est=-667.1+4364*4.5

#The model is
# Y = -667.1+4364*x; units of variables as mentioned in the problem
# The model estimated variance is 5633072
# The regression coefficient is 4364 
# The value of y when x = 4.01 is 16832 kg & the residual is 3107.46 (rounded off to nearest integer)
# The mean mass of crop harvested from 4.5 hectares is 18971 kg (rounded off to nearest integer)

########End of Data 2######

#------------------------------------------------------------------------------------------------------------------#

######Start of Data 3######

#Data 3 = d3

# Fit a model

m3=lm(y_~x1_+x2_+x3_,data=d3)
sigma_sq3=(summary(m3)$sigma)**2
a1=62
a2=27
a3=86
y_cap=-39.92+0.7156*a1+1.30*a2+(-0.16)*a3

#The Regression model is given by for data 3 = d3 (with stackflow as response variable Y) as 
#Y = -39.92+0.7156*x1+1.30*x2-0.16*x3
# The estimated model variance is 10.52 (upto 2 decimal places)
# The fitted value of Y (stackflow) for given x1=62, x2=27 & x3=86 is given by 25.79 (upto 2 decimal places)

######End of Data 3#######

#-------------------------------------------------------------------------------------------------------------------#

######Start of Data 4#######

#Data 4 = d4

# Fit a model

m4=lm(y11~x11+x21,data=d4)
sigma_sq4=(summary(m4)$sigma)**2

#m4-> Y = 3494.5+577.84*x1-78.13*x2

se_b0=as.numeric(summary(m4)$coef[,2][1])
se_b1=as.numeric(summary(m4)$coef[,2][2])
se_b2=as.numeric(summary(m4)$coef[,2][3])

#x2=40, x1=4

x1=4; x2=40

Y_pred = 3494.5+577.84*x1-78.13*x2

#The regression model to predict the number of Internet users is: (all corrected to 2 decimal places)
#Y = 3494.5+577.84*x1-78.13*x2
#The standard error of intercept of the model is 873.72
#The standard error of the regression coefficient of x1 (no. of yrs since research) of the model is 122.35
#The standard error of the regression coefficient of x2 (age) of the model is 16.75
#The predicted value of Y for x1=4 & x2=40 is 2680.66

######End of Data 4######

#-----------------------------------------------------------------------------------------------------------------#

(Answer)